Hendrixapeals:

If 4+2=6 then why isn't 4*2=6

b3lla2025:

cause 4 times two equals 8

supie:

\(*= \times\) So 4 two times \(4*2=4+4\) \(4+4=8\\4*2=8\)

jhonyy9:

4×2 so 4 time 2 is 2+2+2+2 what = 8